### v1

@@ Abstract
We propose a minimal extension of the Standard Model (SM) that addresses both the smallness of neutrino masses and the dark matter (DM) puzzle via the inverse seesaw mechanism and an axion portal fermionic DM. This model generates light neutrino masses without requiring high energy scales, enhancing its testability in future collider experiments. An axion-like particle (ALP) connects the SM and DM sectors, yielding a distinct phenomenology and a positive contribution to the muon anomalous magnetic moment ($g-2$) anomaly. Our analysis shows that the model is consistent with constraints from neutrino oscillations, DM relic density, and indirect detection, as well as providing a solution to the $g-2$ anomaly. The TeV-scale scalar sector and ALP interactions suggest potential signals in collider experiments and astrophysical observations. This work offers a unified framework to address neutrino masses and DM, with implications for particle physics and cosmology.
@@ 1 Introduction
One of the most profound issues with the SM is its prediction of massless neutrinos. However, experimental evidence from Super-Kamiokande Fukuda et al. (1998), SNO Ahmad et al. (2002), KamLAND Eguchi et al. (2003), Daya Bay An et al. (2024), RENOAhn et al. (2012), CHOOZ Apollonio et al. (1999), K2K Ahn et al. (2003), T2KAbe et al. (2011), and LSND Aguilar et al. (2001) confirms that neutrinos undergo flavor oscillations, and hence are massive. Traditional seesaw mechanisms have been proposed to explain the smallness of neutrino masses by introducing heavy right-handed Majorana neutrinos, which generate light neutrino masses suppressed by the heavy Majorana scale Weinberg (1979); Mohapatra and Senjanovic (1980); Schechter and Valle (1980); King (2004); Mohapatra and Valle (1986); Dev and Pilaftsis (2012), commonly referred to as the seesaw scale. However, the traditional seesaw mechanism usually requires a seesaw scale around $M\sim 10^{12}$ GeV, which greatly reduces the possibility to probe the model experimentally. On the other hand, the inverse seesaw model King and Luhn (2013); Deppisch et al. (2015); Gavela et al. (2009); Ibarra et al. (2010); Deppisch and Valle (2005); Dev and Mohapatra (2010); Das and Okada (2013) offers a more testable neutrino mass generation mechanism. In this scenario, sterile neutrinos with small, yet technically natural, lepton number-violation (LNV) couplings are introduced. These small LNV couplings help to lower the seesaw scale down to within reach of colliders. This approach has gained traction as a realistic and testable framework, with studies exploring its implications on lepton flavor violation (LFV) and collider phenomenology Langacker (1981); Davidson et al. (2008); Altarelli and Feruglio (2010); Abada and Lucente (2014); Deppisch et al. (2015); Alekhin et al. (2016); Hisano et al. (1996); Kuno and Okada (2001); Beacham et al. (2020).
@@
On the other hand, the existence of DM has been firmly established through astrophysical observations, such as galaxy rotation curves Navas et al. (2024); Bertone et al. (2005) and cosmic microwave background measurements Aghanim et al. (2020); Spergel et al. (2003); Smoot et al. (1992). Various extensions of the SM have been proposed to accommodate DM, see Ref. Feng (2010); Arcadi et al. (2018); Bertone and Tait (2018); Cirelli et al. (2024); Bozorgnia et al. (2024); Balazs et al. (2024) for recent reviews. One of the most studied scenario is the Higgs portal DM Patt and Wilczek (2006), thanks to its simplicity and predictive nature. In this framework, the Higgs boson serves as a mediator between the DM and the SM sector. Phenomenology of the Higgs portal DM has been studied in Ref. Djouadi et al. (2012); Bishara et al. (2016); Arcadi et al. (2020). Due to its minimal structure, the Higgs portal model cannot accommodate neutrino masses without a further extension.
@@
An alternative and compelling approach is the axion portal model, which introduces an axion or axion-like particle (ALP) as a mediator between the DM and the SM sectors. Initially proposed to address the strong CP problem in quantum chromodynamics, axion has gained interest in DM studies due to their weak interactions and stability Dine et al. (1981); Peccei and Quinn (1977); Weinberg (1978). ALP portal models offers unique signatures in indirect detection experiment, cosmic microwave background distortions, and low energy precision measurements, which can differentiate them from Higgs portal models Nomura and Thaler (2009); Freytsis et al. (2010); Bauer et al. (2023); Zhevlakov et al. (2022). It is also interesting to note that ALP can be connected to neutrino mass generation models as well Carvajal et al. (2015).
Another intriguing anomaly is the muon anomalous magnetic moment, $g-2$, where current experimental measurements at Fermilab Abi et al. (2021); Aguillard et al. (2023) and previous measurements at Brookhaven National Laboratory Bennett et al. (2006) exceed the SM prediction Aoyama et al. (2020). This discrepancy, standing at about $5\sigma$, suggests possible contributions from physics beyond the SM.  [^1] The recent lattice simulation Borsanyi et al. (2021) and experimental results Ignatov et al. (2024a, b) have called into question the validity of the SM prediction. Various models, including those with extended Higgs sectors Chen et al. (2021); Primulando et al. (2023); Iguro et al. (2023), sterile neutrinos Pinheiro et al. (2021); Ashry et al. (2023), or new gauge bosons Gninenko and Krasnikov (2001, 2018), attempt to account for this deviation by introducing additional contributions to $g-2$. In particular, models involving dark sectors or ALPs have been investigated for their potential to explain the muon $g-2$ anomaly Buen-Abad et al. (2021); Ganguly et al. (2022); Janish and Ramani (2020); deNiverville et al. (2018).
@@
In this work, we build upon these developments by proposing a minimal SM extension that combines the inverse seesaw mechanism for neutrino mass generation and the ALP portal DM. Our model draws from the previous study that explored Higgs portals in the DM-neutrino mass model Pongkitivanichkul et al. (2019), aiming instead to utilize the ALP portal Fitzpatrick et al. (2023); Armando et al. (2024); Ghosh et al. (2024); Gutiérrez et al. (2021); Di Meglio (2021); Gola et al. (2022); Deniverville et al. (2021); Kirpichnikov et al. (2020); deNiverville and Lee (2019); deNiverville et al. (2018); Kaneta et al. (2017), which offers new dynamics and observables. In this work, we also attempt to explain the muon $g-2$ anomaly within the ALP portal model, which place a non-trivial constraint on the model parameter space. This approach not only addresses the neutrino mass and DM issues within a unified framework but also proposes testable signatures that could guide future experimental searches.
@@ 2 The model
In this work, we propose an extension to the SM with an electroweak (EW) singlet complex scalar ($\Phi$), two EW singlet right-handed heavy neutrinos ($N^{i}_{R}$), two EW singlet left-handed heavy neutrinos ($S^{I}_{L}$), and a pair of fermionic dark matter ($\chi_{L}$ and $\chi_{R}$).  [^2] Scenarios with one $N_{R}$ and one $S_{L}$, or one $N_{R}$ and two $S_{L}$ lead to two vanishing light neutrino masses. Scenario with two $N_{R}$ and one $S_{L}$ is compatible with neutrino oscillation data. However, it would require more free parameters than the scenario presented in this work. In addition, the $U(1)_{PQ}$ global symmetry is imposed to stabilise the DM sector, and is assumed to be broken at a scale higher than the EW symmetry breaking scale. The relevant particle contents and their charge assignment under $SU(2)_{L}\times U(1)_{Y}\times U(1)_{PQ}$ is summarised in Tab. 1.
@@ 2.1 The scalar sector
$H=\frac{1}{\sqrt{2}}\begin{pmatrix}\sqrt{2}G^{+}\\ v_{h}+h^{\prime}+iG^{0}\end{pmatrix},\qquad\Phi=\frac{1}{\sqrt{2}}(v_{\phi}+\phi^{\prime})e^{ia/f_{a}},$ (2)
where $G^{+}$ and $G^{0}$ are the would-be Goldstone boson. $v_{h}$, $v_{\phi}$ and $f_{a}$ are the vacuum expectation values (vev) of $H$ and $\Phi$ and the ALP decay constant, respectively. Since $U(1)_{PQ}$ is assumed to be broken at scale higher than the EW scale, we have $v_{\phi},f_{a}>v_{h}$. Note that we employ the non-linear parametrization for $\Phi$ which makes the ALP, $a$, appears explicitly as the phase. The fields $h^{\prime}$ and $\phi^{\prime}$ mix to form the mass eigenstates
@@
Thus, the scalar potential can be described by three free parameters: $v_{\phi}$, $m_{\phi}$ and $\sin{\theta}$. We note that $\sin\theta$ is constrained by the 125 GeV Higgs boson measurement. The latest results from the ATLAS and CMS collaborations place an upper bound on $\sin\theta\leq 0.12$ Sirunyan et al. (2021); Aad et al. (2022).
The ALP potential, on the other hand, is generated from the non-perturbative effect, which breaks $U(1)_{PQ}$ into a shift symmetry, $a\to a+f_{a}$. This shift symmetry dictates that the ALP potential takes the form of the instanton cosine potential Adams et al. (2022)
@@
$f_{a}=v_{\phi}.$ (8)
From now on we will use $f_{a}$ in place of $v_{\phi}$. The ALP potential also generates the ALP mass
@@
Below the $U(1)_{PQ}$ symmetry breaking scale, $f_{a}$, the interaction of ALP with other particles is given by the effective operators
$\mathcal{L}_{\rm ALP}=$ $\frac{C_{\psi\psi}}{f_{a}}(\partial_{\mu}a)\overline{\psi}\gamma^{\mu}\gamma_{5}\psi+\frac{g_{s}^{2}}{16\pi^{2}}C_{gg}\frac{a}{f_{a}}G_{\mu\nu}^{a}\tilde{G}^{a\mu\nu}$
$+\frac{g^{2}}{16\pi^{2}}C_{WW}\frac{a}{f_{a}}W_{\mu\nu}^{i}\tilde{W}^{i\mu\nu}+\frac{g^{\prime 2}}{16\pi^{2}}C_{BB}\frac{a}{f_{a}}B_{\mu\nu}\tilde{B}^{\mu\nu},$ (10)
where $\psi$ is the SM fermion with the ALP coupling $C_{\psi\psi}$. The couplings $g_{s}$, $g$, and $g^{\prime}$ are the strong, weak, and hypercharge gauge couplings respectively, and $C_{VV}$ is the gauge-ALP coupling. The dual field strength tensor is defined as $\tilde{F}_{\mu\nu}=\frac{1}{2}\epsilon_{\mu\nu\alpha\beta}F^{\alpha\beta}$. The fields $B$ and $W_{3}$ mix to become the photon and the $Z$-boson via
$\begin{pmatrix}Z\\ A\end{pmatrix}=\begin{pmatrix}c_{W}&-s_{W}\\ s_{W}&c_{W}&\end{pmatrix}\begin{pmatrix}W^{3}\\ B\end{pmatrix},$ (11)
where $s_{W}$ and $c_{W}$ are the sine and cosine of the Weinberg angle respectively. We also define $W^{\pm}_{\mu}=(W^{1}_{\mu}\mp iW^{2}_{\mu})/\sqrt{2}$. Then, the gauge-ALP interaction can be recast into the form
$\mathcal{L}_{\rm gauge-ALP}$ $=C_{\gamma\gamma}\frac{\alpha}{4\pi}\frac{a}{f_{a}}F\tilde{F}+C_{ZZ}\frac{\alpha}{4\pi s_{W}^{2}c_{W}^{2}}\frac{a}{f_{a}}Z\tilde{Z}$
$+\Bigg{(}C_{WW}\frac{\alpha}{4\pi s_{W}^{2}}\frac{a}{f_{a}}W\tilde{W}+C_{\gamma Z}\frac{\alpha}{4\pi s_{W}c_{W}}\frac{a}{f_{a}}A\tilde{Z}+{\rm h.c.}\Bigg{)},$ (12)
where we introduce a shorthand notation $F\tilde{F}=F_{\mu\nu}\tilde{F}^{\mu\nu}$, $\alpha=e^{2}/4\pi$ is the fine structure constant, and
$C_{\gamma\gamma}=C_{WW}+C_{BB},\qquad C_{\gamma Z}=c_{W}^{2}C_{WW}-s_{W}^{2}C_{BB},\quad C_{ZZ}=c_{W}^{4}C_{WW}+s_{W}^{4}C_{BB}.$ (13)
@@ 2.2 The neutrino sector
$\mathcal{L}_{Yuk}\supset-Y^{\alpha i}\overline{L}^{\alpha}\tilde{H}N_{R}^{i}-Y^{\prime Ii}\Phi\overline{S_{L}}^{I}N_{R}^{i}+\mathrm{h.c.},$ (14)
@@
where we define $m_{D}^{\alpha i}=Y^{\alpha i}v_{h}/\sqrt{2}$ and $M_{N}^{Ij}=Y^{\prime Ij}f_{a}/\sqrt{2}$. The neutrino mass terms can be written in a compact form
$-\mathcal{L}_{\rm mass}^{\nu}=$ $\frac{1}{2}\overline{\psi^{c}_{R}}\mathcal{M}_{\nu}\psi_{R}+\mathrm{h.c.},$ (17)
@@
$\psi_{R}=\begin{pmatrix}\nu_{L}^{c}\\ N_{R}\\ S_{L}^{c}\end{pmatrix},\qquad\mathcal{M}_{\nu}=\begin{pmatrix}\mathbf{0}&m_{D}&\mathbf{0}\\ m^{T}_{D}&\mu_{R}&M^{T}_{N}\\ \mathbf{0}&M_{N}&\mu_{S}\end{pmatrix}.$ (18)
The neutrino mass matrix $\mathcal{M}_{\nu}$ can be brought to a block diagonal form by Dev and Pilaftsis (2012)
@@ 2.3 Dark matter sector
The $U(1)_{PQ}$ symmetry breaking generates a DM mass, $m_{\chi}=g_{\chi}f_{a}/\sqrt{2}$. Below the $U(1)_{PQ}$ breaking scale, DM interactions are given by
$\mathcal{L}_{DM}\supset$ $-\frac{m_{\chi}}{f_{a}}\sin{\theta}\;h\overline{\chi}\chi-\frac{m_{\chi}}{f_{a}}\cos{\theta}\;\phi\overline{\chi}\chi-i\frac{m_{\chi}}{f_{a}}\;a\overline{\chi}\gamma_{5}\chi$
$-i\frac{g_{\chi}\cos{\theta}}{\sqrt{2}f_{a}}a\phi\overline{\chi}\chi-i\frac{g_{\chi}\sin{\theta}}{\sqrt{2}f_{a}}ah\overline{\chi}\chi,$ (22)
@@
From the above interaction, one see that the $h$, $\phi$ and $a$ can serve as a portal between the DM sector and the SM sector. However, the $h$ and $\phi$ portal to SM fermions and gauge bosons are suppressed by $\sin^{2}2\theta\lesssim 0.06$ due to the mixing in the neutral scalar sector. The ALP portal, on the other hand, does not suffer from this suppression. For simplicity, we will assume $\theta=0$ in this work. Thus, DM phenomenology is mediated by the ALP portal.

### v2

@@ Abstract
We propose a minimal extension of the Standard Model (SM) that addresses both the smallness of neutrino masses and the dark matter (DM) puzzle via the inverse seesaw mechanism and an axion portal fermionic DM. This model generates light neutrino masses without requiring high energy scales, enhancing its testability in future collider experiments. An axion-like particle (ALP) connects the SM and DM sectors, yielding a distinct phenomenology. Our analysis shows that the model is consistent with constraints from neutrino oscillations and DM relic density as well as satisfying the current measurement on muon $g-2$. This work offers a unified framework to address neutrino masses and DM, with implications for particle physics and cosmology.
@@ 1 Introduction
One of the most profound issues with the SM is its prediction of massless neutrinos. However, experimental evidence from Super-Kamiokande Super-Kamiokande:1998kpq , SNO SNO:2002tuh , KamLAND KamLAND:2002uet , Daya Bay DayaBay:2024hrv , RENORENO:2012mkc , CHOOZ CHOOZ:1999hei , K2K K2K:2002icj , T2KT2K:2011ypd , and LSND LSND:2001aii confirms that neutrinos undergo flavor oscillations, and hence are massive. Traditional seesaw mechanisms have been proposed to explain the smallness of neutrino masses by introducing heavy right-handed Majorana neutrinos, which generate light neutrino masses suppressed by the heavy Majorana scale Weinberg:1979sa ; Mohapatra:1979ia ; Schechter:1980gr ; King:2003jb ; Mohapatra:1986bd ; Dev:2012sg , commonly referred to as the seesaw scale. However, the traditional seesaw mechanism usually requires a seesaw scale around $M\sim 10^{12}$ GeV, which greatly reduces the possibility to probe the model experimentally. On the other hand, the inverse seesaw model King:2013eh ; Deppisch:2015qwa ; Gavela:2009cd ; Ibarra:2010xw ; Deppisch:2004fa ; Dev:2009aw ; Das:2012ze ; Bonilla:2023egs ; Bharadwaj:2024crt offers a more testable neutrino mass generation mechanism. In this scenario, sterile neutrinos with small, yet technically natural, lepton number-violation (LNV) couplings are introduced. These small LNV couplings help to lower the seesaw scale down to within reach of colliders. This approach has gained traction as a realistic and testable framework, with studies exploring its implications on lepton flavor violation (LFV) and collider phenomenology Langacker:1980js ; Davidson:2008bu ; Altarelli:2010gt ; Abada:2014vea ; Deppisch:2015qwa ; Alekhin:2015byh ; Hisano:1995cp ; Kuno:1999jp ; Beacham:2019nyx .
@@
On the other hand, the existence of DM has been firmly established through astrophysical observations, such as galaxy rotation curves ParticleDataGroup:2024cfk ; Bertone:2004pz and cosmic microwave background measurements Planck:2018vyg ; WMAP:2003elm ; COBE:1992syq . Various extensions of the SM have been proposed to accommodate DM, see Ref. Feng:2010gw ; Arcadi:2017kky ; Bertone:2018krk ; Cirelli:2024ssz ; Bozorgnia:2024pwk ; Balazs:2024uyj for recent reviews. One of the most studied scenario is the Higgs portal DM Patt:2006fw , thanks to its simplicity and predictive nature. In this framework, the Higgs boson serves as a mediator between the DM and the SM sector. Phenomenology of the Higgs portal DM has been studied in Ref. Djouadi:2011aa ; Bishara:2015cha ; Arcadi:2019lka . Due to its minimal structure, the Higgs portal model cannot accommodate neutrino masses without a further extension.
@@
An alternative and compelling approach is the axion portal model, which introduces an axion or axion-like particle (ALP) as a mediator between the DM and the SM sectors. Initially proposed to address the strong CP problem in quantum chromodynamics, axion has gained interest in DM studies due to their weak interactions and stability Dine:1981rt ; Peccei:1977hh ; Weinberg:1977ma . ALP portal models offers unique signatures in indirect detection experiment, cosmic microwave background distortions, and low energy precision measurements, which can differentiate them from Higgs portal models Nomura:2008ru ; Freytsis:2009ct ; Bauer:2022rwf ; Zhevlakov:2022vio . It is also interesting to note that ALP can be connected to neutrino mass generation models as well Carvajal:2015dxa .
Another historically interesting observable is the muon anomalous magnetic moment, $g-2$, whose precise measurements at Fermilab Muong-2:2021ojo ; Muong-2:2023cdq and previously at Brookhaven Muong-2:2006rrc have long appeared to deviate from earlier SM predictions Aoyama:2020ynm . This tension, which once stood at nearly $5\sigma$, motivated significant theoretical and experimental efforts. However, a recent comprehensive update by the Muon $g-2$ Theory Initiative Aliberti:2025beg , incorporating improved lattice QCD evaluations and addressing tensions in $e^{+}e^{-}$ data, has led to a revised SM prediction that is in agreement with the experimental average within uncertaintiesMuong-2:2025xyk . While the anomaly is no longer statistically significant, the muon magnetic moment remains a sensitive probe for new physics.
@@
In this work, we build upon these developments by proposing a minimal SM extension that combines the inverse seesaw mechanism for neutrino mass generation and the ALP portal DM. Our model draws from the previous study that explored Higgs portals in the DM-neutrino mass model Pongkitivanichkul:2019cvm , aiming instead to utilize the ALP portal Fitzpatrick:2023xks ; Armando:2023zwz ; Ghosh:2023tyz ; Gutierrez:2021gol ; DiMeglio:2021iec ; Gola:2021abm ; Deniverville:2020rbv ; Kirpichnikov:2020tcf ; deNiverville:2019xsx ; deNiverville:2018hrc ; Kaneta:2017wfh , which offers new dynamics and observables. In this work, we also apply the recent constraint Aliberti:2025beg on muon $g-2$ to the ALP portal model, which place a non-trivial constraint on the model parameter space. This approach not only addresses the neutrino mass and DM issues within a unified framework but also proposes testable signatures that could guide future experimental searches.
@@ 2 The model
In this work, we propose an extension to the SM with an electroweak (EW) singlet complex scalar ($\Phi$), two EW singlet right-handed heavy neutrinos ($N^{i}_{R}$), two EW singlet left-handed heavy neutrinos ($S^{I}_{L}$), a pair of fermionic dark matter ($\chi_{L}$ and $\chi_{R}$).  [^1] Scenarios with one $N_{R}$ and one $S_{L}$, or one $N_{R}$ and two $S_{L}$ lead to two vanishing light neutrino masses. Scenario with two $N_{R}$ and one $S_{L}$ is compatible with neutrino oscillation data. However, it would require more free parameters than the scenario presented in this work. In addition, the chiral $U(1)_{PQ}$ global symmetry is imposed to stabilise the DM sector, and is assumed to be broken at a scale higher than the EW symmetry breaking scale. The relevant particle contents and their charge assignment under $SU(2)_{L}\times U(1)_{Y}\times U(1)_{PQ}$ is summarised in Tab. 1. With this charge assignment, the model does not generate the axion coupling to any SM gauge boson through the triangle diagram of fermions. There is only axion-fermion coupling which is naturally generated by the PQ transformation on the kinetic term of fermions.
@@ 2.1 The scalar sector
$H=\frac{1}{\sqrt{2}}\begin{pmatrix}\sqrt{2}G^{+}\\ v_{h}+h^{\prime}+iG^{0}\end{pmatrix},\qquad\Phi=\frac{1}{\sqrt{2}}(v_{\phi}+\phi^{\prime})e^{ia/v_{\phi}},$ (2)
where $G^{+}$ and $G^{0}$ are the would-be Goldstone bosons. $v_{h}$ and $v_{\phi}$ are the vacuum expectation values (vev) of $H$ and $\Phi$ respectively and $f_{a}$ is the ALP decay constant. Since $U(1)_{PQ}$ is assumed to be broken at scale higher than the EW scale, we have $v_{\phi},f_{a}>v_{h}$. Note that we employ the non-linear parametrization for $\Phi$ which makes the ALP, $a$, appears explicitly as the phase. The fields $h^{\prime}$ and $\phi^{\prime}$ mix to form the mass eigenstates
@@
Thus, the scalar potential can be described by three free parameters: $v_{\phi}$, $m_{\phi}$ and $\sin{\theta}$. We note that $\sin\theta$ is constrained by the 125 GeV Higgs boson measurement. The latest results from the ATLAS and CMS collaborations place an upper bound on $\sin\theta\leq 0.15$ CMS:2022dwd ; ATLAS:2022vkf .
The ALP potential, on the other hand, is generated from the non-perturbative effect, which breaks $U(1)_{PQ}$ into a shift symmetry, $a\to a+f_{a}$. This shift symmetry dictates that the ALP potential takes the form of the instanton cosine potential Adams:2022pbo
@@
$f_{a}=X_{\Phi}v_{\phi}.$ (8)
The ALP potential also generates the ALP mass
@@ 2.2 The neutrino sector
$\mathcal{L}_{Yuk}\supset-Y^{\alpha i}\overline{L}_{L}^{\alpha}\tilde{H}N_{R}^{i}-Y^{\prime Ii}\Phi\overline{S_{L}}^{I}N_{R}^{i}+\mathrm{h.c.},$ (10)
@@
where we define $m_{D}^{\alpha i}=Y^{\alpha i}v_{h}/\sqrt{2}$ and $M_{N}^{Ij}=Y^{\prime Ij}v_{\phi}/\sqrt{2}$. The neutrino mass terms can be written in a compact form
$-\mathcal{L}_{\rm mass}^{\nu}=$ $\frac{1}{2}\overline{\mathcal{N}^{c}_{R}}\mathcal{M}_{\nu}\mathcal{N}_{R}+\mathrm{h.c.},$ (13)
@@
$\mathcal{N}_{R}=\begin{pmatrix}\nu_{L}^{c}\\ N_{R}\\ S_{L}^{c}\end{pmatrix},\qquad\mathcal{M}_{\nu}=\begin{pmatrix}\mathbf{0}&m_{D}&\mathbf{0}\\ m^{T}_{D}&\mu_{R}&M^{T}_{N}\\ \mathbf{0}&M_{N}&\mu_{S}\end{pmatrix}.$ (14)
The neutrino mass matrix $\mathcal{M}_{\nu}$ can be brought to a block diagonal form by Dev:2012sg
@@ 2.3 Dark matter sector
The $U(1)_{PQ}$ symmetry breaking generates a DM mass, $m_{\chi}=g_{\chi}v_{\phi}/\sqrt{2}$. Below the $U(1)_{PQ}$ breaking scale, DM interactions are given by
$\mathcal{L}_{DM}\supset$ $-\frac{m_{\chi}}{v_{\phi}}\sin{\theta}\;h\overline{\chi}\chi-\frac{m_{\chi}}{v_{\phi}}\cos{\theta}\;\phi\overline{\chi}\chi-i\frac{m_{\chi}}{v_{\phi}}\;a\overline{\chi}\gamma_{5}\chi$
$-i\frac{g_{\chi}\cos{\theta}}{\sqrt{2}v_{\phi}}a\phi\overline{\chi}\chi-i\frac{g_{\chi}\sin{\theta}}{\sqrt{2}v_{\phi}}ah\overline{\chi}\chi,$ (18)
@@
From the above interaction, one see that the $h$, $\phi$ and $a$ can serve as a portal between the DM sector and the SM sector. However, the $h$ and $\phi$ portal to SM fermions and gauge bosons are suppressed by $\sin^{2}2\theta\lesssim 0.08$ due to the mixing in the neutral scalar sector. The ALP portal, on the other hand, does not suffer from this suppression. For simplicity, we will assume $\theta=0$ in this work. Thus, DM phenomenology is mediated by the ALP portal.
2.4 The axion interactions
The interaction term of axion-fermion and axion-photon can be derived by demanding that the Lagrangian of the model is invariant under the following PQ transformation
$f_{R}$ $\mapsto$ $e^{-iX_{f}\;a/f_{a}}f_{R},\quad\psi_{R}\mapsto e^{-iX_{\psi}\;a/f_{a}}\psi_{R},\quad\varphi\mapsto e^{-iX_{\varphi}a/f_{a}}\varphi,$ (19)
where $f$ is the SM fermion, $\psi=\{\chi,N\}$ and $\varphi=\{H,\Phi\}$. The $X_{f}$ and $X_{\varphi}$ are PQ-charges of fermions and scalars respectively. According to table 1, the PQ transformation (19) on the kinetic term of SM fermions leads to axion-SM fermion derivative interactions in the form of
$\mathcal{L}_{\rm kin}$ $\supset$ $\overline{f}i\gamma^{\mu}\partial_{\mu}f\to\overline{f}i\gamma^{\mu}\partial_{\mu}f+C_{ff}\frac{\partial_{\mu}a}{f_{a}}\overline{f}\gamma^{\mu}\gamma_{5}f,$ (20)
where $C_{ff}=X_{f}=\pm X_{\varphi}$ where $+$ is for up-type quarks and -- is for charged lepton and down-type quarks.